determine the slope of the line perpendicular to the given line.
3x-2y=5

Answer by ewatrrr(24785)


Hi,
determine the slope of the line perpendicular to the line: 3x-2y=5
*Note:The standard point-slope form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
3x-2y=5 Solving the given line for y to put into the point-slope format
3x -5 = 2y
(3/2)x - (5/2) = y Slope of this line is (3/2)
Perpendicular lines have slopes that are negative reciprocals of one another
(-2/3) is the slope of the line perpendicular to the given line